If the correlation coefficient between two variables is 0.5 and the standard deviation of the variables is 5 and 10, respectively, then the sample covariance is approximately ...

a.

25

b.

50

c.

250

d.

Impossible to determine with any level of accuracy.
Answer #1

Answer
Correlation coefficient =covariance /(product of sandard deviations)
0.5=covariance /(5*10)
covariance =50*0.5
covariance =25
the covariance is 25

option 1
Answer #2

ANSWER :


OPTION a : 25


[ Correlation coefficient, rxy =  Covariancexy / (SDx * SDy)

=> 0.5 = Covariancexy / (5 * 10)

=> covariancexy = 0.5 * 5 * 10 = 25 ]


answered by: Tulsiram Garg